Budget Preparation

Westport Company is preparing its master budget for May. Use the estimates provided to determine the amounts necessary for each of the following requirements. (Estimates may be related to more than one requirement.)

a. What should total sales revenue be if territories E and W estimate sales of 50,000 and 100,000 units, respectively, and the unit selling price is $50?
$Answer

b. If the beginning finished goods inventory is an estimated 6,000 units and the desired ending inventory is 9,500 units, how many units should be produced?
Answer units

c. What dollar amount of materials should be purchased at $4 per pound if each unit of product requires 2.5 pounds and beginning and ending materials inventories should be 12,000 and 13,500 pounds, respectively.
$Answer

d. How much direct labor cost should be incurred if each unit produced requires 0.5 hours at an hourly rate of $19?
$Answer

e. How much manufacturing overhead should be incurred if fixed manufacturing overhead is $45,000 and variable manufacturing overhead is $3.50 per direct labor hour?
$Answer
Answer #1

a) Total Sales revenue = (50000+100000)*50 = $7500000

b) Unit to be produced = 150000+9500-6000 = 153500 Units

c) Raw material purchase = (153500*2.5)+13500-12000 = 385250

Raw material purchase cost = 385250*4 = $1541000

d) Direct labor cost = 153500*.50*19 = $1458250

e) Total manufacturing overhead = (153500*.50*3.5)+45000 = 313625